OpenStudy (jh99):

how are you supposed to convert y=(x+2)(x-3) into standard and vertex form? I got y=x^2-x-6 for standard form, but I don't think this is right because I can't convert it to vertex form.

satellite73 (satellite73):

looks right to me

OpenStudy (jh99):

how do you convert that to vertex form? It doesn't work

OpenStudy (loser66):

why not?

OpenStudy (loser66):

\(x^2-x-6= x^2-x+(1/4)-(1/4)-6\) \(=(x-1/4)^2 -25/4\)

OpenStudy (loser66):

ooops!! mistake \((x-(1/2))^2-(25/4)\)

OpenStudy (loser66):

got it?

OpenStudy (jh99):

Ohh thank you, it was queer because there was a fraction in it. That messed me up! Thanks for helping :)

OpenStudy (loser66):

np